OpenStudy (anonymous):

How do you solve rational equations? For example, (k+4)/4 + (k-1)/4=k+4/(4k). How do you check for extraneous solutions?

OpenStudy (anonymous):

what is the LCM of k+4/4 and k-1/4 ???

OpenStudy (anonymous):

Well, I was just wondering how to check for extraneous solutions. I've gimplified the equation to 8k^2+8k-16=0, but I'm not sure if that's right.

OpenStudy (anonymous):

(k+4)/4 +(k-1)/4 =(k+4)/4k (k+4+k-1)/4 = (k+4)/4k (2k+3)/4 = (k+4)/4k (2k+3)k = (k+4) 2k^2+3k=k+4 2k^2+2k-4=0 k^2+k-2=0

OpenStudy (anonymous):

that looks correct

OpenStudy (anonymous):

now try to factor k^2+k-2=0

OpenStudy (anonymous):

OR u can use the direct formula

OpenStudy (anonymous):

(k-1)(k+1)=0?

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

k^2+k-2=0 k^2+2k-k-2=0 k(k+2)-1(k+2)=0 (k-1)(k+2)=0

OpenStudy (anonymous):

does it help?

OpenStudy (anonymous):

Yup, much thanks. However, what does it mean when it says "check for extraneous solutions"?

OpenStudy (anonymous):

may be for imaginery solutions

OpenStudy (anonymous):

How can you tell when you need to check?

OpenStudy (anonymous):

if ax^2+bx+c=0 where b^2-4ac<0

OpenStudy (anonymous):

but I am also not sure what extraneous solution means

OpenStudy (anonymous):

Oh the discriminant

OpenStudy (anonymous):

yes